OpenStudy (anonymous):

i am having trouble understanding how to figure out how to do interpreting data for 8th grade pre alg.

OpenStudy (anonymous):

question??

OpenStudy (anonymous):

a surey of 200 randomly selected consumersfinds that 75 consumers perfer brand a dtergent.predict how mant consumers in a town of 3000 perfer brand detergent.

OpenStudy (anonymous):

75 yes of 200 total How many yes (x?) out of 3000

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

You remember similar triangles in geometry? 5/7=x/10? Like that.

OpenStudy (anonymous):

no dont remeber...........never been in geomerty

OpenStudy (anonymous):

You have 75 consumers out of 200, that can be written as 75/200 You want to know, let's call it x, out of 3000 that can be written as x/3000 The two are similar so you can set them equal and solve for x x/3000=75/200

OpenStudy (anonymous):

ohhh okay thank you.